OpenStudy (ccastudentgrade10):

FAN&MEDAL Find the length of the hypotenuse of a triangle is side a = sqrt15 and side b = 2sqrt15. Do not use a calculator. help? I have to do three of these

OpenStudy (ccastudentgrade10):

*if side a

OpenStudy (mathstudent55):

Since the triangle has a hypotenuse, what kind of triangle is it?

OpenStudy (ccastudentgrade10):

its a right triangle @mathstudent55

OpenStudy (ccastudentgrade10):

i just don't know how to solve it without a calculator

OpenStudy (mathstudent55):

Since you know it's a right triangle, you can use the Pythagorean theorem. \(a^2 + b^2 = c^2\) Use the two given sides as a and b, and solve for c.

OpenStudy (mathstudent55):

\(a^2 + b^2 = c^2\) First, copy the equation above. Now rewrite the equation, but replace a with \(\sqrt{15}\), and replace b with \(2\sqrt{15} \). What do you have now?

OpenStudy (ccastudentgrade10):

would it be 15 + 30 = c^2?

OpenStudy (ccastudentgrade10):

so then sqrt45 would be the answer

OpenStudy (mathstudent55):

No.

OpenStudy (mathstudent55):

First just replace a and b by their values. Then calculate.

OpenStudy (mathstudent55):

\(a^2 + b^2 = c^2\) \((\sqrt{15})^2 + (2\sqrt{15})^2= c^2\) You see how the substitution into the formula is done?

OpenStudy (ccastudentgrade10):

oh okay. but the sqrt15^2 just becomes 15, right?

OpenStudy (mathstudent55):

When you square \(\sqrt {15} \), you do get 15. When you square \(2 \sqrt {15} \), you must also square the 2.

OpenStudy (mathstudent55):

\((\sqrt{15})^2 + (2\sqrt{15})^2= c^2\) \(15 + 2^2 (\sqrt{15})^2 = c^2\) \(15 + 4 \times 15 = c^2\) The 2 is also squared.

OpenStudy (ccastudentgrade10):

OH okay

OpenStudy (ccastudentgrade10):

so 15 + 60 = c^2 75 = c^2 75^2 = c

OpenStudy (mathstudent55):

No. You are correct up to here: \(75 = c^2\) Now you take square roots of both sides. We can also switch sides first to have the variable on the left. \(c^2 = 75\) Take square roots. We will only use the positive square root of 75 since we are dealing with the length of a side of a triangle which cannot be negative. \(c = \sqrt {75} \) Ok so far?

OpenStudy (mathstudent55):

There is one final step we can do. We can simplify the square root. We need to factor out a perfect square number out of 75.

OpenStudy (ccastudentgrade10):

okay

OpenStudy (ccastudentgrade10):

my teacher says 75^2 is fine, i can just leave the answer squared

OpenStudy (mathstudent55):

\(c = \sqrt {75} = \sqrt{25 \times 3} = \sqrt{25}\sqrt{3} = 5 \sqrt 3\)

OpenStudy (ccastudentgrade10):

thank you!!

OpenStudy (mathstudent55):

75^2 is completely incorrect. You can leave the answer as sqrt(75), or \(\sqrt {75} \), but not as \(75^2\)

OpenStudy (mathstudent55):

You're welcome.